OpenStudy (anonymous):

The base of a triangle is 7 cm greater than the height. The area is 39cm^2. Find the ht and length of the base.

OpenStudy (watchmath):

Base =B. Height=H. We have B= H +7. Area = (1/2)B*H = (1/2)(H+7)H = 39. Multiply by 2 we have \(H^2+7H=78\) \(H^2+7H-78=0\) \(H+13)(H-6)=0\) Since H is positive, then H=6. It follows that B=6+7=13.

OpenStudy (anonymous):

if base 13 then do i * 13 by 6 to get the length

OpenStudy (watchmath):

The problem wants us to find the height and the (length) of the base. The answer are H=6 and B=13.

OpenStudy (anonymous):

o, i c